OpenStudy (anonymous):

a bacteria culture initially contains 100 cells and grows at a proportional rate to its size. after an hr the population has increased to 420 a find the expression of bacteria after three hrs b). find the rate of growth after three hrs @Mathematics

OpenStudy (anonymous):

let no., at any instant be N. gn\[dN/dt \alpha N\] dN/dt=kN........(a) integrating,\[\int\limits_{?}^{?}dN/N=k \int\limits_{?}^{?}dt\] lnN=kt+c.........(1) when ,t=0 & N=100 ln100=c c=2.....(2) lnN=kt+2 when , t=1&N=420 ln420=k+2 {2.30300 * log(420) = 6.04134312} k=4.04.....(3) [lnN=4.04t+c] is the required eqn. from (a) when t=3, dN/dt=4.04*N.......(b) but ,lnN=4.04*3+2 2.303*logN=12.12+2 logN=14.12/2.303 logN= 6.1311333 N=1353000........(c) from (a)&(b) dN/dt=4.04*1353000 dN/dt=5 466 120per day